a normal population has sd 2.1.The standard error of the sampling distribution of the sample mean with sample size 900 is



1
Expert's answer
2021-07-19T05:41:52-0400
"se=\dfrac{\sigma}{\sqrt{n}}=\dfrac{2.1}{\sqrt{900}}=\dfrac{2.1}{30}=0.07"